OpenStudy (anonymous):

What is a reciprocal? How would I do it on a fraction? All answers get medals!

OpenStudy (mertsj):

Tip the number upside down. That is the reciprocal. \[\frac{3}{2} \]

OpenStudy (mertsj):

\[\frac{2}{3}\]

OpenStudy (mertsj):

\[1\frac{2}{3}=\frac{5}{3}\]

OpenStudy (mertsj):

\[\frac{3}{5}\]